A 600-g solid sphere has a diameter of 10 cm rotates at 600 revolutions per minute about an

axis through its center. Determine the kinetic energy of the system.


1
Expert's answer
2021-03-16T11:38:37-0400

Mass M=0.600 kg

Radius is R=0.05 m

Angular speed is "\omega =600 rev/min = 62.83 rad/s"

The kinetic energy of the system is

"K= \frac{1}{2}I \omega^2= \frac{1}{2}( \frac{2}{5} MR^2) \omega^2"

"=\frac{1}{5} MR^2 \omega^2=\frac{1}{5} (0.60)(0.05)^2(62.83)^2=1.184 J"